OpenStudy (anonymous):

A segment has endpoints at (2, -15) and (8, 4). What is the y-coordinate of the midpoint of that segment? Enter your answer as a decimal.

OpenStudy (e.mccormick):

The midpoint is half the distance. So, you need to look at how much the x and y values are changing, and only change half as much.

OpenStudy (e.mccormick):

Another way of doing this is with what they call the midpoint formula. In \((x,y)\) it makes the point: \(\left (\dfrac{x_1+x_2}{2},\dfrac{y_1+y_2}{2}\right )\)

OpenStudy (anonymous):

(5, -11) ???

OpenStudy (e.mccormick):

The x is 5, but the y is not -11. you forgot to divide that by 2.

OpenStudy (e.mccormick):

And since it is asking for the y value, that is the one you must get right.